# Revisiting the propagation of highly-energetic gamma rays in the Galaxy

### v4

@@ I Introduction
The past decade has been a fervent and vivid era for gamma-ray astronomy field thanks to the activity of space- and ground-based experiments. Among the former are the Fermi-LAT (Large Area Telescope) [1] and AGILE (Astro-Rilevatore Gamma a Immagini Leggero[2] detectors, amid the latter are MAGIC (Major Atmospheric Gamma Imaging Cherenkov) [3] and H. E. S. S. (High Energy Stereoscopic System) [4] telescopes, LHAASO (Large High Altitude Air Shower Observatory) [5] and HAWC (High-Altitude Water Cherenkov) [6] observatories. Moreover, the approaching advent of CTA (Cherenkov Telescope Array) [7], ASTRI (Astrofisica con Specchi a Tecnologia Replicante Italiana) Mini-Array [8] and SWGO (Southern Wide-field Gamma-ray Observatory) [9] observatories will enable the exploration of the gamma-ray sky with unprecedented performances [10]. In this context, recent detection of galactic gamma rays with energies ranging from several hundreds of TeV up to few PeVs [11, 12], in some cases showing extended morphologies at the highest energy bands [13, 14], pose new theoretical challenges in the understanding of the sources.
Emissions of gamma rays with energies beyond tens of TeV are likely produced by galactic PeVatrons, efficient accelerators of cosmic rays up to several PeVs [15, 16, 11, 17, 18, 19]. Both hadronic and/or leptonic mechanisms have been proposed to explain the consequent gamma-ray emission from PeVatrons. In the former, gamma rays are generated from decays of neutral pions, secondaries of hadronic processes that involve cosmic rays accelerated by PeVatrons [20, 21, 22]. Leptonic scenarios explain the production of gamma rays with energy larger than 100 TeV in young pulsars placed in radiation-dominated regions [23, 24, 25].
@@
Apart from production mechanisms, gamma rays propagate over astrophysical distances before eventually reaching our observatories. During their travel, they might interact with background photons permeating astrophysical spaces, producing electron-positron pairs, to which from now on we refer simply as electrons. Such electrons are then deflected by astrophysical magnetic fields and, in turn, they might scatter through inverse Compton (IC) background photons up to several TeVs. These processes trigger electromagnetic (EM) cascades over astrophysical distances [^1] Throughout the text, even with a low number of occurring interactions we refer to the phenomena as EM cascade.. In this picture, prompt gamma-ray fluxes could be distorted during their travel towards detectors on Earth. Specifically, their energy spectra could present absorption features and arrival direction maps with spreading counts due to “deflected” gamma rays. The latter might lead to the formation of haloes, i.e. extended gamma-ray counts surrounding the emission cores. Moreover, in this picture, the deflection of electrons in the EM cascade produce delays in the arrival times of gamma rays. These effects have been largely studied in the context of gamma rays propagating over intergalactic lengths [48], leading to constraints on intergalactic magnetic field properties [49, 50, 51, 52]. Propagation effects are fundamental not only in the understanding of source production mechanisms, but also for peculiar indirect searches for axion-like particles [53, 54, 55, 56, 57, 58], dark matter candidates. The goal of this paper is to explore this scenario for gamma rays travelling over galactic length scales.
@@
To achieve our goals, it is crucial to characterise the galactic environment. Besides the cosmic microwave background (CMB) and cosmic radio background (CRB), galactic space is permeated by the interstellar radiation field (ISRF). In intergalactic space, the equivalent of the ISRF is the extragalactic background light (EBL) [59, 60, 61, 62, 63, 64, 65]. Previous works investigated gamma-ray galactic absorption due to pair production on ISRF and CMB by computing the optical depth [66, 67, 68, 69, 70, 71]. They found that gamma-ray fluxes from sources located in the galactic center start to be attenuated at energies above $\sim~{}10$ TeV. Moreover, our Galaxy has a peculiar magnetic field structure, product of thermal electrons pervading interstellar space, large-scale plasma bubbles and small-scale turbulent features due to outflows, then enhanced by hydrodynamic turbulences. Galactic magnetic field (GMF) models are few and carry uncertainties, given the complexity and variety of the measurements involved [72, 73, 74, 75].
@@
This work aims to investigate theoretical scenarios of gamma rays, with energy between 100 GeV and 100 PeV, propagating in our Galaxy, using dedicated Monte Carlo (MC) simulation tools. It is organised as follows: in Sec. II the simulation settings are described together with gamma-ray propagation theory and galactic environmental properties. In Sec. III, the observables at Earth position from the simulated gamma-ray sources are examined. Chiefly, spectral and spatial observables are related to the specific source position in the Galaxy, because of background photons and magnetic field spatial models. Sec. IV outlines the conclusions of this work and open questions to address in future follow-up studies.
@@ II.1 Interactions and background photon fields
  • triplet pair production [84]: $e+\gamma_{\text{bkg}}\to e+e^{-}+e^{+}$
@@
– It is the background photon field due to the integrated history of galaxy and star emissions. These lower energy photons usually dominate the attenuation of gamma rays with energies larger than $10^{18}$ eV. The model adopted in this work is the one reported in Ref. [95], a refinement of a previous study [103]. As shown in Fig. 2, IC off CRB photons starts to be relevant for hundreds-of-TeV electrons, dominating at energies beyond 100 PeV.
@@ II.2 Galactic magnetic field model
In reproducing galactic environments, the GMF model employed in this paper is the one reported in Ref. [75] (JF12+ hereafter), an improved version of Ref. [72]. It comprises three components: the large-scale regular field, small-scale turbulent fields, with coherence length of 100 pc or smaller, and the striated random field, aligned over larger scales, but with varying strength and sign. The colour maps of Fig. 3 represent the magnetic field intensity at the galactic plane for the three contributions combined. The regular component is, in turn, made up of three items: the disc, the toroidal halo, and the out-of-plane component.
@@ II.3 Spherical observer & gamma-ray sources
The ten simulated sources are placed around the Galaxy as the white circles in Fig. 3, all of them pointing towards Earth. There are two sets of sources specular with respect to the x-axis, hereafter $S_{\lessgtr}$ because of their $\lessgtr 0$ y-coordinate. Each contains four sources, with a line of sight that goes through the inner regions of our Galaxy. For these two sets, the ISRF domains implemented are delimited by red lines in Fig. 3. The other two sources are 7 kpc (alias $S_{1}$) and 5.6 kpc ($S_{2}$) far from Earth: the related ISRF regions are bounded by blue lines in Fig. 3. The lines of sight cross, respectively, one and two galactic spiral arms, seen in the GMF of Fig. 3.
@@ IV Conclusion and prospects
The aim of this work is to investigate the role of secondary effects in the propagation of gamma rays in our Galaxy, determined by the environmental characteristics, i.e. background photons and magnetic fields. These drive the development of EM cascades, triggered by gamma rays with energies larger than tens of TeV.
@@
It is important to stress the limitations of the results here reported. The ISRF implemented in the simulations is an approximation of the model: a future work might require the incorporation of a space-dependent photon field in the code framework. We are actively working on implementing this in the CRPropa code. Besides code concerns, this work assumes specific models for both the ISRF and the GMF. It would be worthwhile to investigate various ISRF and GMF models present in literature, with all the possible updates and scenarios. For instance, a very recent revision of the GMF model proposing eight coherent component models is Ref. [121]. Actually, it would impact on the ratio $B_{\text{turb}}/B_{\text{reg}}$, thus consequently on the arrival direction maps.
Overall, the significant point would be to compare what learnt from simulations, in terms of gamma-ray absorption and “deflection”, with current and future gamma-ray experiments performance, to properly infer the impact of propagation effects on observed galactic gamma-ray fluxes. This will result in improved characterisations of PeVatrons and galactic center intrinsic emission models, as well as in more robust constraints from dark matter searches through gamma rays.
@@ Acknowledgements.
The simulation data performed and employed in this work are accessible for downloading at the data repository of the DAMASCO group website. The files contain also the EM cascade electrons reaching the observer, despite they are not accounted in our analysis. We encourage those interested in the data to reach out the authors for discussions about data structuring and analysis.
@@ Appendix B Observer size
The observer of the simulations, positioned at the nominal Earth position, is spherical in shape. In our work, the sphere size is geometrically related to the angular aperture of the intrinsic emission cone of the source, given that the jet axis points towards the center of the observer sphere. The geometrical radius is $R=D\cdot\tan\phi$, with $\phi$ denoting half-aperture of the prompt emission cone and $D$ the observer-source distance. However, we add a tiny angle $\delta$ to the half-aperture $\phi$ since, if a gamma ray originates with an initial direction in the outer layers of the cone, the observer may detect minute deflections. Therefore the observer radii chosen in our simulations are $R_{\text{obs}}=D\cdot\tan(\phi+\delta)$, as depicted in Fig. 16.
@@ Appendix C Arrival direction maps
This appendix contains the spatial count maps for the ten sources simulated in this work, as discussed in Sec. III.2. The count maps of $S_{\lessgtr}$ sources are in Fig. 22 and Fig. 27, respectively for the closest and farthest sources. The ones corresponding to $S_{1}$ and $S_{2}$ are shown in Fig. 29 and Fig. 29, respectively. The amount of injected gamma rays in the simulation of each source is $10^{7}$. To note that the counts on the colorbars are not effective since, in our simulations, the thinning procedure is applied to sample the secondary particles in the cascade [77].

### v5

@@ I Introduction
The past decade has been a fervent and vivid era for gamma-ray astronomy thanks to the activity of space- and ground-based experiments. Among the former are the Fermi-LAT (Large Area Telescope) [1] and AGILE (Astro-Rilevatore Gamma a Immagini Leggero[2] detectors, amid the latter are MAGIC (Major Atmospheric Gamma Imaging Cherenkov) [3] and H. E. S. S. (High Energy Stereoscopic System) [4] telescopes, LHAASO (Large High Altitude Air Shower Observatory) [5] and HAWC (High-Altitude Water Cherenkov) [6] observatories. Moreover, the approaching dawn of CTAO (Cherenkov Telescope Array Observatory) [7], ASTRI (Astrofisica con Specchi a Tecnologia Replicante Italiana) Mini-Array [8] and SWGO (Southern Wide-field Gamma-ray Observatory) [9] observatories will enable the exploration of the gamma-ray sky with unprecedented performances [10]. In this context, recent detection of galactic gamma rays with energies ranging from several hundreds of TeV up to few PeVs [11, 12], in some cases showing extended morphologies at the highest energy bands [13, 14], pose new theoretical challenges in the understanding of the sources.
Emissions of gamma rays with energies beyond tens of TeV are likely produced by galactic PeVatrons, efficient accelerators of cosmic rays up to several PeVs [15, 16, 17, 18, 11, 19, 20, 21]. Both hadronic and/or leptonic mechanisms have been proposed to explain the consequent gamma-ray emission from PeVatrons. In the former, gamma rays are generated from decays of neutral pions, secondaries of hadronic processes that involve cosmic rays accelerated by PeVatrons [22, 23, 24]. Leptonic scenarios explain the production of gamma rays with energy larger than 100 TeV in young pulsars placed in radiation-dominated regions [25, 26, 27].
@@
Apart from production mechanisms, gamma rays propagate over astrophysical distances before eventually reaching our observatories. During their travel, they might interact with background photons permeating astrophysical spaces, producing electron-positron pairs, to which from now on we refer simply as electrons. Such electrons are then deflected by astrophysical magnetic fields and, in turn, they might scatter through inverse Compton (IC) background photons up to several TeVs. These processes trigger electromagnetic (EM) cascades over astrophysical distances [^1] Throughout the text, even with a low number of occurring interactions we refer to the phenomena as EM cascade.. In this picture, prompt gamma-ray fluxes could be distorted during their travel towards detectors on Earth. Specifically, their energy spectra could present absorption features and arrival direction maps with spreading counts due to “deflected” gamma rays. The latter might lead to the formation of haloes, i.e. extended gamma-ray counts surrounding the emission cores. Moreover, in this picture, the deflection of electrons in the EM cascade produce delays in the arrival times of gamma rays. These effects have been largely studied in the context of gamma rays propagating over intergalactic lengths [50], leading to constraints on intergalactic magnetic field properties [51, 52, 53, 54]. Propagation effects are fundamental not only in the understanding of source production mechanisms and intergalactic spaces, but also for peculiar indirect searches for axion-like particles [55, 56, 57, 58, 59, 60], dark matter candidates.
@@
The goal of this paper is to explore the influence of propagation effects on gamma rays travelling over galactic length scales, after leaving their production region. To achieve this, it is crucial to characterise the galactic environment. Besides the cosmic microwave background (CMB) and cosmic radio background (CRB), galactic space is permeated by the interstellar radiation field (ISRF). In intergalactic space, the equivalent of the ISRF is the extragalactic background light (EBL) [61, 62, 63, 64, 65, 66, 67]. Previous works investigated gamma-ray galactic absorption due to pair production on ISRF and CMB by computing the optical depth [68, 69, 70, 71, 72, 73]. They found that gamma-ray fluxes from sources located in the galactic center start to be attenuated at energies above $\sim~{}10$ TeV. Moreover, our Galaxy has a peculiar magnetic field structure, product of thermal electrons pervading interstellar space, large-scale plasma bubbles and small-scale turbulent features due to outflows, then enhanced by hydrodynamic turbulences. Galactic magnetic field (GMF) models are few and carry uncertainties, given the complexity and variety of the measurements involved [74, 75, 76, 77].
@@
This work aims to investigate theoretical scenarios of gamma rays, with energy between 100 GeV and 100 PeV, propagating in our Galaxy, using dedicated Monte Carlo (MC) simulation tools. Unlike the previously cited works, these tools also enable us to trace the three-dimensional development of EM cascades, reconstructing accordingly the galactic environment. It is organised as follows: in Sec. II the simulation settings are described together with gamma-ray propagation theory and galactic environmental properties. In Sec. III, the observables at Earth position from the simulated gamma-ray sources are examined. Chiefly, spectral and spatial observables are related to the specific source position in the Galaxy, because of background photons and magnetic field spatial models. Sec. IV outlines the conclusions of this work and open questions to address in future follow-up studies.
@@ II.1 Interactions and background photon fields
  • triplet pair production [86, 87, 88, 89]: $e+\gamma_{\text{bkg}}\to e+e^{-}+e^{+}$
@@
– It is the background photon field due to the integrated history of galaxy and star emissions. These lower energy photons usually dominate the attenuation of gamma rays with energies larger than $10^{18}$ eV. The model adopted in this work is the one reported in Ref. [100], a refinement of a previous study [108]. The novelty of Ref. [100] lies in the more accurate treatment of the radio processes occurring in star-forming and radio galaxies, as well as their morphology and evolution. As shown in Fig. 2, IC off CRB photons starts to be relevant for hundreds-of-TeV electrons, dominating at energies beyond 100 PeV.
@@ II.2 Galactic magnetic field model
In reproducing galactic environments, the GMF model employed in this paper is the one reported in Ref. [77] (JF12+ hereafter), an implemented version of Ref. [74] – the former only improves the out-of-plane GMF and ensures the magnetic flux conservation. It comprises three components: the large-scale regular field, small-scale turbulent fields, with coherence length of 100 pc or smaller, and the striated random field, aligned over larger scales, but with varying strength and sign. The colour maps of Fig. 3 represent the magnetic field intensity at the galactic plane for the three contributions combined. The regular component is, in turn, made up of three items: the disc, the toroidal halo, and the out-of-plane component.
@@ II.3 Spherical observer & gamma-ray sources
The ten simulated sources are placed around the Galaxy as the white arrows in Fig. 3, all of them pointing towards Earth. There are two sets of sources mirror-symmetric with respect to the x-axis, hereafter $S_{\lessgtr}$ because of their $\lessgtr 0$ y-coordinate. Each contains four sources, with a line of sight that goes through the inner regions of our Galaxy. For these two sets, the ISRF domains implemented are delimited by red lines in Fig. 3. The other two sources are 7 kpc (alias $S_{1}$) and 5.6 kpc ($S_{2}$) far from Earth: the related ISRF regions are bounded by blue lines in Fig. 3. The lines of sight cross, respectively, one and two galactic spiral arms, seen in the GMF of Fig. 3. Interestingly, the distance and line of sight – far from the galactic center – of the $S_{1}$ source resemble the location of 1LHAASO J2002+3244u, one of the multi-TeV gamma-ray sources in the first LHAASO catalog [110].
@@ IV Conclusion and prospects
The aim of this work is to investigate the role of propagation effects on gamma rays traveling through our Galaxy, determined by the environmental characteristics, i.e. background photons and magnetic fields. These drive the development of EM cascades, triggered by gamma rays with energies larger than tens of TeV.
@@
It is important to stress the limitations of the results here reported. The ISRF implemented in the simulations is an approximation of the model: a future work might require the incorporation of a space-dependent photon field in the code framework. We are actively working on implementing this in the CRPropa code. It will be designed for detailed studies of observed gamma-ray sources all around our Galaxy, serving as the natural extension of this work. Besides code concerns, this work assumes specific models for both the ISRF and the GMF. It would be worthwhile to investigate various ISRF and GMF models present in literature, with all the possible updates and scenarios. For instance, a very recent revision of the GMF model, proposing eight coherent component models, is presented in Ref. [126]. Actually, it would impact on the ratio $B_{\text{turb}}/B_{\text{reg}}$, thus on the arrival direction maps. A follow-up interesting work could involve the detailed modelisation of gamma-ray sources or smaller size, i.e. $\lesssim 100\;\text{pc}$, regions in our Galaxy in the code, e.g. OB super-bubbles whose environments [127, 128] are efficient cosmic-ray factories [129, 130].
Overall, the significant point would be to compare what learnt from simulations, in terms of gamma-ray absorption and “deflection”, with current and future gamma-ray experiments performance, to properly infer the impact of propagation effects on observed galactic gamma-ray fluxes. This will result in improved characterisations of PeVatrons and galactic centre intrinsic emission models, as well as in more robust constraints from dark matter searches through gamma rays.
@@ Acknowledgements.
The simulation data performed and employed in this work are accessible for downloading at the data repository of the DAMASCO group website [131]. The files contain also the EM cascade electrons reaching the observer, despite they are not accounted in our analysis. We encourage those interested in the data to reach out the authors for discussions about data structuring and analysis.
@@ Appendix B Observer size
The observer of the simulations, positioned at the nominal Earth position, is spherical in shape. In our work, the sphere size is geometrically related to the angular aperture of the intrinsic emission cone of the source, given that the jet axis points towards the centre of the observer sphere. The geometrical radius is $R=D\cdot\tan\phi$, with $\phi$ denoting half-aperture of the prompt emission cone and $D$ the observer-source distance. However, we add a tiny angle $\delta$ to the half-aperture $\phi$ since, if a gamma ray originates with an initial direction in the outer layers of the cone, the observer may detect minute deflections. Therefore the observer radii chosen in our simulations are $R_{\text{obs}}=D\cdot\tan(\phi+\delta)$, as depicted in Fig. 16.
@@ Appendix C Arrival direction maps
This appendix contains the spatial count maps for the ten sources simulated in this work, as discussed in Sec. III.2. The count maps of $S_{\lessgtr}$ sources are in Fig. 22 and Fig. 27, respectively for the closest and farthest sources. The ones corresponding to $S_{1}$ and $S_{2}$ are shown in Fig. 29 and Fig. 29, respectively. The amount of injected gamma rays in the simulation of each source is $10^{7}$. To note that the values on the colorbars are not actual counts since, in our simulations, the thinning procedure is applied to sample the secondary particles in the cascade [79].